OpenStudy (anonymous):

The formula for a geometric sequence is an=2(4)^n-1. What is the third term of the sequence? A.24 B.32 C.16

OpenStudy (anonymous):

Third term of n = 3 2*(4)^(3 -1) = 2* (4)^ 2 = 2 * 16 = 32

OpenStudy (anonymous):

So the correct answer is B